Belinda is twice as old as Rose now. Seven years from now, the sum of their ages will be 44 years. How old will Rose be in 7 years?

Answer by TimothyLamb(4379)

x = belinda age now
y = rose age now
---
x = 2y
(x + 7) + (y + 7) = 44
---
x = 2y
x + y = 30
---
put the system of linear equations into standard form
---
x - 2y = 0
x + y = 30
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = belinda age now = 20
y = rose age now = 10
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php